Problem:
Write the equation of the line L satisfying the given geometric conditions.
L has y-intercept (0, 2) and is perpendicular to the line with equation
2x-3y=6

Answer by stanbon(75887)

Write the equation of the line L satisfying the given geometric conditions.
L has y-intercept (0, 2) and is perpendicular to the line with equation
2x-3y=6
----------
The given equation can be written y=(2x-6)/3 = (2/3)x-2
Its slope is 2/3
------------
The slope of any line perpendicular to it has slope -3/2.
If that perpendicular line also passes thru (0,2) its
equation is y=(-3/2)x+2 because (0,2) is the y-intercept.
===========
cheers,
Stan H.